1/1

N. N. 32,

A spotless roze is growing

tradukita de N. N. 33

1. A Spotless Rose is growing, Sprung from a tender root, Of ancient seers’ foreshowing, Of Jesse promised fruit; Its fairest bud unfolds to light Amid the cold, cold winter, And in the dark midnight.

2. The Rose which I am singing, Whereof Isaiah said, Is from its sweet root springing In Mary, purest Maid; Through God’s great love and might The Blessed Babe she bare us In a cold, cold winter’s night.

Traduko de la Germana poemo “Es ist ein Ros entsprungen” de N. N. 32 en la Anglan de N. N. 33.

Arg-618-1248 (2012-10-19 08:19:20)

Tiu ĉi traduko troviĝas en http: // ingeb. org/ spiritua/ esistein. html .

http://www.poezio.net/version?poem-id=618&version-id=1248